How do I solve the following equation by the square root method?
(3y+6)^2=36

Answer by checkley77(12844)

(3y+6)^2=36
Take the sqrt of both sides.
3y+6=6
3y=6-6
3y=0
y=0/6
y=0 ans.
Proof:
(3*0+6)^2=36
6^2=36
36=36